OpenStudy (anonymous):

PLEASEE Find the vertex of the parabola formed by the following quadratic function: y = -x2 - 6x - 8 (-3, 1) (3, -35) (6, -80) (-6, -8)

OpenStudy (anonymous):

write equation in form of(y-k)=(x-h) ^2 +c where (h,k) is vertex

OpenStudy (anonymous):

so it would b the 3rd one?

OpenStudy (anonymous):

y = -x2 - 6x - 8 y-(-8)=-x2 - 6x -9+9 n proceed